OpenStudy (anonymous):

Please help, will medal 3 _______ 2x + y

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

nice question

Nnesha (nnesha):

but what u have to do with that question ?????

OpenStudy (anonymous):

simplify it

OpenStudy (anonymous):

That's all it said "simplify the complex fraction"

OpenStudy (anonymous):

It's \[\frac{ 3 }{ 2x+y }\]

OpenStudy (anonymous):

It's completely simplified??

OpenStudy (squirrels):

You can't simplify it anymore. Are you sure they don't want you to find x or y?

OpenStudy (anonymous):

OMG I wrote it wrong!! I'm so sorry just a second

Nnesha (nnesha):

^_^

OpenStudy (anonymous):

\[\frac{ 3 }{ \frac{ 2 }{ x } +y}\]

OpenStudy (anonymous):

That's what needs to be simplified

OpenStudy (anonymous):

\[\frac{ 3x }{ xy+2 }\]

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

multiply it all by x?

OpenStudy (anonymous):

Is this multiple choice or extended response.

OpenStudy (anonymous):

extended response

OpenStudy (anonymous):

Factor the top and the bottom to simplify the answer. Remember: in this problem, change the numerator to a single fraction before you start to simplify. The second term in the numerator will be multiplied by x/x to create the needed denominator.

OpenStudy (anonymous):

Oh! ok that makes sense, so is that the end answer? the one you posted a second ago?

OpenStudy (anonymous):

Yuppers

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

Yes'm xD